OpenStudy (anonymous):

Which polynomial represents a perfect square polynomial? Select all that apply. 1.x^2+8x+16 2.4x^2+10x+25 3.9x^2+48x+64 4.16x^2+4x+1

OpenStudy (mimi_x3):

well lets try to perfect square each one .... A quadratic equation has the following form \(ax^2+bx+c\) and when we perfect square it we have \((ax+\sqrt{c})^2\) and \(b=2*a*\sqrt{c}\) \[x^2+8x+16\] So a in this case a=1 and c=16 and\(\sqrt{16}=4\) and b=8 \[(x+4)^2\] Lets see if \( b=2*a*\sqrt{c}\) so \( 2*1*4=2*4=8\) So the first option is a perfect square